Suppose the production function is Y = AK0.3N0.7.Suppose in 2010, K = 1000, N = 100, and Y = 199.5.

In 2020, capital, labor, and output have doubled, soK = 2000, N = 200, and Y = 399.

 

(a.)        By what percentage did productivity grow from2010 to 2020?

(b.)        If output had risen to 798 instead of 399, and capitaland labor doubled, by what percentage wouldproductivity have grown from 2010 to 2020?



1
Expert's answer
2021-03-14T19:04:50-0400

a) Productivity Improvement = (399 - 199.5) = 199.5

= (199.5 / 199.5) x 100 = 100%

productivity grew by 100%


b) Productivity Improvement = (798 - 199.5) = 598.5

(598.5 / 199.5) x 100 = 300

productivity grew by 300%